Start learning →

At least one '6' in 4 throws of a fair die:

A4/6
B2/6
C1 - (5/6)⁴ ≈ 0.518
D2/3
Answer & Solution
Correct answer: C. 1 - (5/6)⁴ ≈ 0.518
Complementary: P(no 6 in 4) = (5/6)⁴ = 625/1296. P(at least one 6) = 1 - 625/1296 = 671/1296 ≈ 0.518.
Practise this question with Quill — GRE →

Keep practising Statistics and Probability

Browse all Statistics and Probability questions for GRE, or explore more Mathematics topics.

Start a Statistics and Probability lesson →